OpenStudy (anonymous):

Solve the equation. Check your answer. 3x – 7 = 5x + 19

OpenStudy (anonymous):

2x=26 so x=13

sam (.sam.):

3x-7=5x+19 Move all terms not containing x to the right-hand side of the equation. -2x-7=19 move all terms not containing x to the righthand side of the equation -2x=26 Divide each term in the equation by -2. x=-13

OpenStudy (anonymous):

wait sorry it is -13

OpenStudy (swag):

x=-13

OpenStudy (anonymous):

3x – 7 = 5x + 19 5x-3x=-7-19 2x=-26 x=-13